Homework problem- Long term Liabilities (Issue and partial Redemption of bonds between interest payment date)

Star Center issued $2,000,000 of 5%, 20-year bonds. These bonds were issued on January 1, 2016 and pay interest annually on each January 1. The bonds yield 3% and was issued at $2,595,107

Required:

(a) Prepare the journal entry to record the issuance of the bonds on January 1, 2016.

   

b) Prepare a bond amortization schedule up to and including January 1, 2020, using the effective-interest method.

c) Assume that on July 1, 2019, Star Center Co. redeems 50% of the bonds at a cost of $1,358,000 plus accrued interest. Prepare the journal entries to record this redemption.
Answer #1
(a) Journal on issue of bond $ $
Cash    25,95,107
Premium on issue of bond           5,95,107
5% Bond        20,00,000
(b)
Year Cash Interest Interest Expenses Premium amortised Unamortized Premium Balance
(a) (b) (c ) (d=c-b) (e=op. bal.- d)
1 Jan., 2016                     -                     -                     -             5,95,107
1 Jan., 2017        1,00,000      1,29,755          29,755           5,65,352
1 Jan., 2018        1,00,000      1,23,268          23,268           5,42,084
1 Jan., 2019        1,00,000      1,17,104          17,104           5,24,980
1 Jan., 2020        1,00,000      1,11,249          11,249           5,13,731
So, premium on bond payable balance is $ 1,41,110
Working Notes:
b) Cash Interest (2000000 x 5%) = $ 1,00,000
c) Interest expenses on issued value
1 Jan., 2017 (2595107*5%) 1,29,755
1 Jan., 2018 ((2595107-129755)*5%) 1,23,268
1 Jan., 2019 ((2595107-129755-123268)*5%) 1,17,104
1 Jan., 2020 ((2595107-129755-123268-117104)*5%) 1,11,249
(c) Journal on redemption of 50% Bond on July 1, 2019
$ $
5% Bond    10,00,000 10,00,000
Accrued Interest          50,000 (2000000*5%)/12*6 = 50,000
Premium on issue of bond      3,08,000 1358000-1000000-50000
Cash        13,58,000